A train of 150 m length is going towards north at a speed of 10 m/s.A bird is flying at 5 m/s parallel to the track towards south.The time taken by the bird to cross the train is   (a)10s (b)15s  (c)30s  (d)12s

Grade:12th pass

5 Answers

Vijay Mukati
8 years ago
The distance which bird has to travel to cross the train is equal to the length of trian. Now since the train and bird are travelling in opposite direction, therefore the speed will be added up. Hence, time = 150m/(10+5) = 10s. Thanks.
Anisha Khan
8 years ago
Thank you,sir.
Brijesh
6 years ago
Relative concept we can apply here , since train and parrot are moving in different directions but it is given that they are moving parallel hence we can use sign convention for both one as negative as south and positive as north direction and we can find the relative velocity of parrot with respect to train and distance is given then we can find time also.Answer will come same as 10 second .
Shaikh faizan
5 years ago
Two trains are moving with equal speed in opposite direction along to parallel railway tracks. if the wind is blowing with the speed u along the tracks so that the relative velocity of the train with respect to the wind are in the ratio 1:2 then the speed of each train must be
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Velocity of bird w.r.t train = 5 – ( – 10) = 15 m/s
Now, length of train = 150 m
Time taken to cross the train by bird = 150/15 = 10 s
 
Thanks and regards,
Kushagra